OpenStudy (anonymous):

Help! Giving Medal A baseball player has a 0.260 batting average. Find the probability that he will have at least two hits in five times at bat.

OpenStudy (kropot72):

The binomial distribution applies. You can solve it by finding the probabilities of 0 and 1 hits and adding these two probabilities. Subtracting the sum from 1 will give the required probability: \[P(0\ hits)=(1.00-0.26)^{5}=you\ can\ calculate\] \[P(1\ hit)=5C1\times0.26\times(0.74)^{4}=you\ can\ calculate\] \[P(at\ least\ 2\ hits)=(1.00-(P(0\ hits)+P(1\ hit)))\]

OpenStudy (anonymous):

@kropot72 Thanks Really Appreciate it =)

OpenStudy (kropot72):

You're welcome :)